# Collaborative Planar Pushing of Polytopic Objects with Multiple Robots in Complex Scenes

### v1

@@ III-C Hierarchical Hybrid Search
Given the above mode generation method, the section presents the hierarchical hybrid search algorithm to generate the hybrid plan as the timed sequence of modes and target states. In particular, it consists of two main steps: (i) a collision-free guiding path is generated for the target given the initial and goal states; (ii) a hybrid search algorithm the that iteratively decomposes the guiding path into arc segments and selects the optimal modes, to minimize a balanced cost of the resulting hybrid plan.
@@ III-C2 Keyframe-guided Hybrid Search
Given the guiding path of the target, a hybrid plan is determined as a sequence of trajectory segment and the associated interaction mode (including contact points and forces), such that the target can be pushed along the guiding path. As shown in Fig. 8, a hybrid search algorithm called keyframe-guided hybrid search (KG-HS) is presented in this section. It iteratively decomposes the guiding path $\overline{\mathbf{S}}$ into segments of arcs between keyframes, such that one pushing mode is sufficient for each segment.
@@ 19             Return the best $\nu^{\star}$;
such that the original arc $\overgroup{\mathbf{s}_{\ell}\mathbf{s}_{\ell+1}}$ is split into two sub-arcs $\overgroup{\mathbf{s}_{\ell}\mathbf{s}_{\widehat{\ell}}}$ and $\overgroup{\mathbf{s}_{\widehat{\ell}}\mathbf{s}_{\ell+1}}$. This step is encapsulated by $\nu^{+}\triangleq\texttt{INS}\big{(}\nu,\ell,(\mathbf{s}_{\widehat{\ell}},\varnothing)\big{)}$, Meanwhile, to account for possible disturbances, a set of close-by states $\widehat{\mathbf{S}}_{\widehat{\ell}}$ are generated by adding perturbations to $\mathbf{s}_{\widehat{\ell}}$, e.g., small rotation and translation. The associated set of new nodes $\{\nu^{+}\}$ are all added to $Q$; (ii) If $\overgroup{\mathbf{s}_{\ell}\mathbf{s}_{\ell+1}}$ is collision-free, a set of feasible modes is generated by solving the associated optimization in (13), denoted by $\mathbf{\Xi}(\overgroup{\mathbf{s}_{\ell}\mathbf{s}_{\ell+1}})$. For each mode $\xi^{\star}\in\mathbf{\Xi}$, a new node is generated by replacing the keyframe $(\mathbf{s}_{\ell},\varnothing)\in\nu^{\star}$ with $(\mathbf{s}_{\ell},\xi^{\star})$, i.e.,
@@ III-D2 Mode Switching
Initially, since the robots are homogeneous, they are assigned to their closest contact points given the first mode $\xi_{0}$. Afterwards, as the segment $\overgroup{\mathbf{s}_{\ell}\mathbf{s}_{\ell+1}}$ is traversed, the fleets switch from mode $\xi_{\ell}$ to $\xi_{\ell+1}$. To reduce switching time and avoid possible collisions, the switching strategy is designed as follows. Let the set of old contact points be $\{\mathbf{c}_{n}\}$ for mode $\xi_{\ell}$ and the new set $\{\mathbf{c}^{+}_{n}\}$ for mode $\xi_{\ell+1}$. First, the periphery of the object is transformed into a circle such that the relative ordering of the contact points is preserved. Second, an diameter associated with angle $\theta^{\star}\in[0,\,\pi]$ is chosen such that the number of old contact points and the number of new contact points on both each side of the diameter are equal. Consequently, the points in $\{\mathbf{c}^{+}_{n}\}$ are numbered clockwise starting from $\theta=\theta^{\star}$, while the same applies for the robots at $\{\mathbf{c}_{n}\}$. Lastly, each robot is assigned to the new contact point with the same number and a local path is planned for navigation. It can be shown that no collisions between the robots can happen and the length of each path is not longer than half of the perimeter if the minimum distance between the target and obstacles is greater than the robot radius; and all robots move at similar speed, as illustrated in Fig. 2.
@@ III-E3 Generalization
The proposed method can be generalized to different scenarios: (I) Heterogeneous robots. When the robots are heterogeneous with different capabilities, the interaction mode should be defined as $\xi\triangleq(\mathbf{c}_{1},R_{c_{1}})\cdots(\mathbf{c}_{N},R_{c_{N}})$, which assigns specific robot to each contact point $\mathbf{c}_{n}$ based on the associated force constraints. For instance, robots with larger maximum forces are assigned to contact points that require larger forces, while robots with smaller forces are assigned to assistive contact points that provide additional geometric constraints; (II) 3D pushing tasks. If the task is to push a target object with 6D pose in a complex 3D scene, the proposed framework can be adopted with minor modifications, under the condition that the workspace is zero-gravity and and resistant, due to the quasi-static assumption. More specifically, consider a target object of which the state is denoted by $\mathbf{s}(t)\triangleq(\mathbf{x}(t),\mathbf{R}(t))$, where $\mathbf{x}(t)\in\mathbb{R}^{3}$ is the position and $\mathbf{R}(t)\in\mathbb{R}^{3\times 3}$ is the rotation matrix for orientation. Additionally, the robots are modeled as spheres with radius $r$, with the translational velocity $\mathbf{v}_{n}\in\mathbb{R}^{3}$ as control inputs. When moving dynamically, the target experiences a force of resistance that is proportional to its generalized velocity, i.e., $\mathbf{q}_{\texttt{res}}\triangleq-\mathbf{K}_{\texttt{res}}\mathbf{p}\in\mathbb{R}^{6}$ with $\mathbf{K}_{\texttt{res}}$ being a diagonal matrix, and $\mathbf{p}=(\mathbf{v},\bm{\omega})=(v_{x},v_{y},v_{z},\omega_{x},\omega_{y},\omega_{z})\in\mathbb{R}^{6}$ as the generalized velocity. Similar to the 2D case, consider that the robots push the target with velocity $\mathbf{p}(t)=\text{diag}(\mathbf{R}(t),\mathbf{R}(t))\mathbf{p}^{\scriptscriptstyle{}{\mathrm{B}}}$ for a time duration $\bar{t}>0$. Then, the arc transition in (11) can be extended to 3D workspace by:
@@ IV-A2 Results
As is shown in Fig. 2 and Fig. 10, the proposed method successfully completed all tasks in three distinctive scenarios. The intermediate results are also shown including the guiding path, the hybrid plan (with keyframes and modes highlighted in different colors), the resulting trajectory of tracking the hybrid plan via the proposed online NMPC, and the control inputs of all robots during execution. The planning time of three scenarios is $12.3s$, $13.2s$ and $21.1s$, while the execution time is $79.2s$, $117s$ and $64.1s$. The three tasks take $3$, $7$ and $3$ mode switches, which is due to the fact that more rotations are required for the second scenario. The proposed NMPC can effectively track the optimized trajectory, with the average tracking errors being $0.025m$,$0.026m$,and $0.048m$. More metrics are discussed in the comparisons below. Moreover, it is interesting to note that despite the NMPC problem in (22) can be solved in approximately $40ms$, yielding a maximum rate of $25Hz$, higher updating rates for NMPC does not significantly improve control performance. Instead, the tracking control is more sensitive to the update rate of the intermediate goal selected from the hybrid plan. If the goal is chosen too close, the resulting trajectory might have drastic changes of the velocity, while a too far goal might lead to a delayed response to errors. Therefore, a lower rate to update the intermediate goal would allow a more gradual convergence. Lastly, it is worth mentioning that in all three scenarios the online execution of the hybrid plan does not require any replanning, different from the hardware experiments in the sequel.
@@ IV-A5 Generalization
(II) Heterogeneous robots. Three heterogeneous robots are deployed for the same pushing task as in Fig. 11, i.e., the maximum pushing forces $f_{n,\max}$ in (1) are $60N$, $40N$ and $20N$, respectively. The assignment and planning strategy as described in Sec. III-E3 is adopted. The overall task is completed with $4$ modes and $60.8s$. Compared with the nominal case where all robots are identical, the assignments of robots to contact points are significantly different: (i) The robot with the largest force is always assigned to the contact point that requires the largest force (the top-left corner of the object); (ii) the robot with the smallest force is assigned to the assistive contact points (the bottom-right corner). Lastly, if this heterogeneity is neglected, the task can still be completed but with a much longer time of $82.2s$.
@@
(III) 6D pushing in 3D workspace. The extension to pushing tasks in complex 3D workspace is demonstrated here with a L-shape object and different number of robots. The object has a mass of $5kg$ with the resistance coefficient $\mathbf{K}=100$, while the robots have a radius of $0.15m$ and a maximum pushing force of $20N$. As shown in Fig. 17, the workspace has a size of $10m\times 20m\times 5m$ with two walls and a small window. The object starts from $\mathbf{x}_{0}=(3,1,3)$ and its destination is at $\mathbf{x}_{\texttt{G}}=(5,17,3)$. Thus, the guiding path of 6D poses is already complex that fist avoids two walls of different heights from above and below, and then passes through the window. The first row shows the snapshots of $4$ robots performing the task, which requires $10$ modes and a total mission time of $92.2s$. In contrast, the second row shows the results for $6$ robots, which requires $4$ modes and a total mission time of $64.3s$. Thus, similar to the 2D case, a larger fleet often requires less number of mode switches and faster completion.
@@ IV-A6 Comparisons
(II) IAB, which enhances POCB further by introducing a probability of choosing a contact point on the surface, which is inversely correlated with the angle $\Delta\theta_{\texttt{G},\texttt{n}}$. Thus, the robot can prioritize pushing directions that are more aligned towards the target. This adaption is particularly significant for long rectangular objects as in the first scenario.
(III) SDF, which is the same as KG-HS but replaces MDF with single directional feasibility in both mode generation and cost estimation modules;
@@
The tests are repeated $50$ times for each method in all three scenarios, by choosing random initial and goal positions for the target object. Sample trajectories are shown in Fig. 20 for the first scenario. As shown in Table I, seven metrics are compared. It can be seen that our KG-HS algorithm achieves the best performance in all metrics except for the planning time, as both POCB and IAB requires no optimization, while AUS has a simpler function to estimate feasibility. Nonetheless, the KG-HS algorithm has the shortest runtime $57.5s$ by summing up the planning and execution time.
@@ IV-B2 Results
Overall, the experiment results are consistent with the simulations for both free and cluttered environments. Namely, via the proposed method, both tasks are completed successfully within $50s$ and $55s$, which require $2$ and $3$ modes, respectively. The average tracking errors are $0.023m$ and $0.052m$. Compared with simulations, the model uncertainties caused by communication delays and actuation noises or slipping are much more apparent in the hardware experiments. Thus, the replanning mechanism is necessary during online execution, as described in Sec. III-D3. Once the object deviates from the planned trajectory by more than $0.2m$, or remains static for $5s$, the KG-HS algorithm is re-executed to generate a new hybrid plan, which is then tracked by the NMPC controller. As shown in Fig. 23, the system in some cases can not recover from the large deviation from the reference trajectory in the current interaction mode. Thus, a replanning is triggered, after which another mode is adopted and the task is completed. Detailed experiment videos are available in the supplementary material.
@@ V Conclusion
There are some limitations of the proposed approach that are worth mentioning: (i) The quasi-static analyses in Sec. III-A1 are only valid for slow movements of the object and the robots. In high-speed scenarios, the object may experience large instantaneous accelerations, which could require the robotic fleet to react with larger forces and torques that may not be allowed in the current pushing mode. Although this can be partially mitigated by the multi-directional feasibility estimation proposed in (10) and the inertia term to NMPC in (22), a comprehensive solution for the high-speed scenarios remains part of our future work; (ii) The intrinsics of the object such as mass, center of gravity and friction coefficient are all assumed to be known for the evaluation MG-SO in (10). Although slight deviations can be handled by the NMPC in (22) and the re-planning process as described in Sec. III-D3, large uncertainties might require other techniques, e.g., model-free and learning-based methods that adapt pushing strategies online. This is part of our ongoing work; (iii) The workspace is assumed to be feasible in the sense that a guiding path $\overline{\mathbf{S}}$ always exists for the given initial and goal states. Despite of being rather mild, it can be further relaxed, e.g., by allowing movable obstacles or by temporarily pulling out some robots to make room and re-join later, which is part of future work; (iv) Lastly, a centralized perception system is assumed that has a full knowledge of the system state and the environment. A decentralized method for perception and state estimation would further improve the applicability.

### v2

@@ III-C Hierarchical Hybrid Search
Given the above mode generation method, the section presents the hierarchical hybrid search algorithm to generate the hybrid plan as the timed sequence of modes and target states. In particular, it consists of two main steps: (i) a collision-free guiding path is generated for the target given the initial and goal states; (ii) a hybrid search algorithm that iteratively decomposes the guiding path into arc segments and selects the optimal modes, to minimize a balanced cost of the resulting hybrid plan.
@@ III-C2 Keyframe-guided Hybrid Search
Given the guiding path of the target, a hybrid plan is determined as a sequence of trajectory segment and the associated interaction mode (including contact points and forces), such that the target can be pushed along the guiding path. As shown in Fig. 8, a hybrid search algorithm called keyframe-guided hybrid search (KG-HS) is presented in this section. It iteratively decomposes the guiding path $\overline{\mathbf{S}}$ into segments of arcs between keyframes, such that each segment can be tracked by single pushing mode.
@@ 19             Return the best $\nu^{\star}$;
such that the original arc $\overgroup{\mathbf{s}_{\ell}\mathbf{s}_{\ell+1}}$ is split into two sub-arcs $\overgroup{\mathbf{s}_{\ell}\mathbf{s}_{\widehat{\ell}}}$ and $\overgroup{\mathbf{s}_{\widehat{\ell}}\mathbf{s}_{\ell+1}}$. This step is encapsulated by $\nu^{+}\triangleq\texttt{INS}\big{(}\nu,\ell,(\mathbf{s}_{\widehat{\ell}},\varnothing)\big{)}$. Meanwhile, to account for possible disturbances, a set of close-by states $\widehat{\mathbf{S}}_{\widehat{\ell}}$ are generated by adding perturbations to $\mathbf{s}_{\widehat{\ell}}$, e.g., small rotation and translation. The associated set of new nodes $\{\nu^{+}\}$ are all added to $Q$; (ii) If $\overgroup{\mathbf{s}_{\ell}\mathbf{s}_{\ell+1}}$ is collision-free, a set of feasible modes is generated by solving the associated optimization in (13), denoted by $\mathbf{\Xi}(\overgroup{\mathbf{s}_{\ell}\mathbf{s}_{\ell+1}})$. For each mode $\xi^{\star}\in\mathbf{\Xi}$, a new node is generated by replacing the keyframe $(\mathbf{s}_{\ell},\varnothing)\in\nu^{\star}$ with $(\mathbf{s}_{\ell},\xi^{\star})$, i.e.,
@@ III-D2 Mode Switching
Initially, since the robots are homogeneous, they are assigned to their closest contact points given the first mode $\xi_{0}$. Afterwards, as the segment $\overgroup{\mathbf{s}_{\ell}\mathbf{s}_{\ell+1}}$ is traversed, the fleets switch from mode $\xi_{\ell}$ to $\xi_{\ell+1}$. To reduce switching time and avoid possible collisions, the switching strategy is designed as follows. Let the set of old contact points be $\{\mathbf{c}_{n}\}$ for mode $\xi_{\ell}$ and the new set $\{\mathbf{c}^{+}_{n}\}$ for mode $\xi_{\ell+1}$. First, the periphery of the object is transformed into a circle such that the relative ordering of the contact points is preserved. Second, an diameter associated with angle $\theta^{\star}\in[0,\,\pi]$ is chosen such that the number of old contact points and the number of new contact points on both side of the diameter are equal. Consequently, the points in $\{\mathbf{c}^{+}_{n}\}$ are numbered clockwise starting from $\theta=\theta^{\star}$, while the same applies for the robots at $\{\mathbf{c}_{n}\}$. Lastly, each robot is assigned to the new contact point with the same number and a local path is planned for navigation. It can be shown that no collisions between the robots can happen and the length of each path is not longer than half of the perimeter, if the minimum distance between the target and obstacles is greater than the robot radius; and all robots move at similar speed, as illustrated in Fig. 2.
@@ III-E3 Generalization
The proposed method can be generalized to different scenarios: (I) Heterogeneous robots. When the robots are heterogeneous with different capabilities, the interaction mode should be defined as $\xi\triangleq(\mathbf{c}_{1},R_{c_{1}})\cdots(\mathbf{c}_{N},R_{c_{N}})$, which assigns specific robot to each contact point $\mathbf{c}_{n}$ based on the associated force constraints. For instance, robots with larger maximum forces are assigned to contact points that require larger forces, while robots with smaller forces are assigned to assistive contact points that provide additional geometric constraints; (II) 3D pushing tasks. If the task is to push a target object with 6D pose in a complex 3D scene, the proposed framework can be adopted with minor modifications, under the condition that the workspace is zero-gravity and resistant, due to the quasi-static assumption. More specifically, consider a target object of which the state is denoted by $\mathbf{s}(t)\triangleq(\mathbf{x}(t),\mathbf{R}(t))$, where $\mathbf{x}(t)\in\mathbb{R}^{3}$ is the position and $\mathbf{R}(t)\in\mathbb{R}^{3\times 3}$ is the rotation matrix for orientation. Additionally, the robots are modeled as spheres with radius $r$, with the translational velocity $\mathbf{v}_{n}\in\mathbb{R}^{3}$ as control inputs. When moving dynamically, the target experiences a force of resistance that is proportional to its generalized velocity, i.e., $\mathbf{q}_{\texttt{res}}\triangleq-\mathbf{K}_{\texttt{res}}\mathbf{p}\in\mathbb{R}^{6}$ with $\mathbf{K}_{\texttt{res}}$ being a diagonal matrix, and $\mathbf{p}=(\mathbf{v},\bm{\omega})=(v_{x},v_{y},v_{z},\omega_{x},\omega_{y},\omega_{z})\in\mathbb{R}^{6}$ as the generalized velocity. Similar to the 2D case, consider that the robots push the target with velocity $\mathbf{p}(t)=\text{diag}(\mathbf{R}(t),\mathbf{R}(t))\mathbf{p}^{\scriptscriptstyle{}{\mathrm{B}}}$ for a time duration $\bar{t}>0$. Then, the arc transition in (11) can be extended to 3D workspace by:
@@ IV-A2 Results
As is shown in Fig. 2 and Fig. 10, the proposed method successfully completed all tasks in three distinctive scenarios. The intermediate results are also shown including the guiding path, the hybrid plan (with keyframes and modes highlighted in different colors), the resulting trajectory of tracking the hybrid plan via the proposed online NMPC, and the control inputs of all robots during execution. The planning time of three scenarios is $12.3s$, $13.2s$ and $21.1s$, while the execution time is $79.2s$, $117s$ and $64.1s$. The three tasks take $3$, $7$ and $3$ mode switches, which is due to the fact that more rotations are required for the second scenario. The proposed NMPC can effectively track the optimized trajectory, with the average tracking errors being $0.025m$,$0.026m$,and $0.048m$. More metrics are discussed in the comparisons below. Moreover, it is interesting to note that despite the NMPC problem in (22) can be solved in approximately $40ms$, yielding a maximum rate of $25Hz$, higher updating rates for NMPC does not significantly improve control performance. Instead, the tracking control is more sensitive to the update rate of the intermediate goal selected from the hybrid plan. If the goal is chosen too close, the resulting trajectory might have drastic changes of the velocity, while a too far goal might lead to a delayed response to errors. Therefore, a lower rate to update the intermediate goal would allow a more gradual convergence. Lastly, it is worth mentioning that in all three scenarios the online execution of the hybrid plan does not require any replanning, different from the hardware experiments.
@@ IV-A5 Generalization
(II) Heterogeneous robots. Three heterogeneous robots are deployed for the same pushing task as in Fig. 11, i.e., the maximum pushing forces $f_{n,\max}$ in (1) are $60N$, $40N$ and $20N$, respectively. The assignment and planning strategy as described in Sec. III-E3 is adopted, completing the task with $4$ modes and $60.8s$. Compared with the nominal case where all robots are identical, the assignments of robots to contact points are significantly different: (i) The robot with the largest force is always assigned to the contact point that requires the largest force (the top-left corner); (ii) the robot with the smallest force is assigned to the assistive contact points (the bottom-right corner). Lastly, if this heterogeneity is neglected, the task can still be completed but with a much longer time of $82.2s$.
@@
(III) 6D pushing in 3D workspace. The extension to pushing tasks in complex 3D workspace is demonstrated here with a L-shape object and different number of robots. The object has a mass of $5kg$ with the resistance coefficient $\mathbf{K}=100$, while the robots have a radius of $0.15m$ and a maximum pushing force of $20N$. As shown in Fig. 17, the workspace has a size of $10m\times 20m\times 5m$ with two walls and a small window. The first row shows the snapshots of $4$ robots performing the task, which requires $10$ modes and a total mission time of $92.2s$. In contrast, the second row shows the results for $6$ robots, which requires $4$ modes and a total mission time of $64.3s$. Thus, similar to the 2D case, a larger fleet often requires less number of mode switches and faster completion.
@@ IV-A6 Comparisons
(II) IAB, which enhances POCB by introducing a probability of choosing a contact point, which is inversely correlated with the angle $\Delta\theta_{\texttt{G},\texttt{n}}$. Thus, the robot can prioritize pushing directions that are more aligned towards the target, which is particularly significant for long rectangular objects in Fig. 20.
(III) SDF, which replaces MDF in KGHS with single directional feasibility in both mode generation and cost estimation;
@@
The tests are repeated $50$ times for each method in all three scenarios, by choosing random initial and goal positions for the target object. Sample trajectories are shown in Fig. 20 for the first scenario. As shown in Table I, seven metrics are compared. It can be seen that our KG-HS algorithm achieves the best performance in all metrics except for the planning time, as both POCB and IAB requires no optimization, while AUS performs less times of mode generations. Nonetheless, the KG-HS algorithm has the shortest runtime $57.5s$ by summing up the planning and execution time.
@@ IV-B2 Results
Overall, the experiment results are consistent with the simulations for both free and cluttered environments. Namely, via the proposed method, both tasks are completed successfully within $50s$ and $55s$, which require $2$ and $3$ modes, respectively. The average tracking errors are $0.023m$ and $0.052m$. Compared with simulations, the model uncertainties caused by communication delays and actuation noises or slipping are much more apparent in the hardware experiments. Thus, the replanning mechanism is necessary during online execution, as described in Sec. III-D3. Once the object deviates from the planned trajectory by more than $0.2m$, or remains static for $5s$, the KG-HS algorithm is re-executed to generate a new hybrid plan, which is then tracked by the NMPC controller. As shown in Fig. 23, the system in some cases can not recover from the large deviation from the reference trajectory in the current interaction mode. Thus, a replanning is triggered, after which another mode is adopted.
@@ V Conclusion
There are some limitations of the proposed approach that are worth mentioning: (i) The quasi-static analyses in Sec. III-A1 are only valid for slow movements of the object and the robots. In high-speed scenarios, the object may experience large instantaneous accelerations, which could require the robotic fleet to react with larger forces and torques that may not be allowed in the current pushing mode. Although this can be partially mitigated by the multi-directional feasibility estimation proposed in (10) and the inertia term to NMPC in (22), a comprehensive solution for the high-speed scenarios remains part of our future work; (ii) The intrinsics of the object such as mass, center of gravity and friction coefficient are all assumed to be known for the evaluation of MG-SO in (10). Although slight deviations can be handled by the NMPC in (22) and the re-planning process as described in Sec. III-D3, large uncertainties might require other techniques, e.g., model-free and learning-based methods that adapt pushing strategies online. This is part of our ongoing work; (iii) The workspace is assumed to be feasible in the sense that a guiding path $\overline{\mathbf{S}}$ always exists for the given initial and goal states. Despite of being rather mild, it can be further relaxed, e.g., by allowing movable obstacles or by temporarily pulling out some robots to make room and re-join later, which is part of future work; (iv) Lastly, a centralized perception system is assumed that has a full knowledge of the system state and the environment. A decentralized method for perception and state estimation would further improve the applicability.
@@
Acknowledgement
This work was supported by the National Key Research and Development Program of China under grant 2023YFB4706500; the National Natural Science Foundation of China (NSFC) under grants 62203017, U2241214, T2121002; and the Fundamental Research Funds for the central universities.